OpenStudy (anonymous):

The square root of 10/81?

OpenStudy (anonymous):

\[\sqrt{a/b}=\sqrt{a}/\sqrt{b}\]

OpenStudy (anonymous):

\[\sqrt{\frac{10}{81}}=\frac{\sqrt{10}}{\sqrt{81}}=\frac{\sqrt{10}}{9}\]